# Experimental Demonstration of an On-Axis Laser Ranging Interferometer for Future Gravity Missions

### v1

@@ IV.1 Angle calibration
Residual alignment imperfections in the QPR introduce cross-coupling effects in the measurements. As shown in Fig. 9 (c), (d), (e), and (f), jumps in cross-coupling signals are observed when the hexapod reverses its scanning direction, while the cross-coupling signals remain relatively constant between these jumps. These jumps may result from the hysteresis and backlash effects [NewportHXPManual], which produce transient instability of the hexapod’s motion during its settling time. Excluding the observed jumps, a linear fit of the measured cross-coupling DWS and DPS signals with respect to the rotation angle determines the cross-coupling factors ( and ), as presented in Table 1.
@@ IV.2 Beam Pointing
The stability of the beam pointing in an LRI system affects the laser link, and the pointing noise can be coupled into ranging measurements through TTL coupling. To investigate the pointing stability of the on-axis architecture, the hexapod was commanded to perform typical motions derived from GRACE-FO pointing data recorded on January 1st, 2019 [goswami2021analysis], which artificially produce the TX beam pointing errors. The hexapod motion was updated at a frequency of 2 Hz over 15-hour continuous measurements, with the CoM as the nominal pivot point. The beam steering loops were closed during the hexapod motions. The coordinate information, provided from the hexapod’s position sensors, was used to monitor hexapod motion in real time. Accordingly, the resulting amplitude spectral density (ASD) of the hexapod motion is shown in Fig. 10 (a), which is consistent with the result in [goswami2021analysis]. Although several harmonics above 0.1 Hz are present, these frequency components are at the boundary of the GRACE-FO frequency band of 2 mHz to 0.1 Hz [schutze2014laser] and thus do not affect our analysis here.
To evaluate the co-alignment between the RX beam and the TX beam, the DWS signals from QPR1 and QPR2 were recorded. For comparison, the DPS signals from QPR3 were used to characterize the absolute TX beam pointing. Using the relation of Eq. 1, the pointing angles were determined by multiplying the DWS and DPS signals by the inverse of the corresponding conversion matrices detailed in Table 1. The corresponding ASD of the pointing angle is displayed in Fig. 9 (b). In addition to the in-loop DWS measurements from the DWS1, out-of-loop DWS measurements from the DWS2 were performed as well. The horizontal (yaw) and vertical (pitch) co-alignments derived from the in-loop DWS measurements are below 10-10 rad/ at the frequency range from 0.07 mHz to 0.5 Hz, which indicates that the control loops work properly. For frequencies between 1 mHz and 0.1 Hz, both direction co-alignments derived from the out-of-loop DWS measurements are lower than 10 µrad/, which fulfills the pointing requirement of GRACE-FO.
@@ IV.3 Polarization
In the beam steering loop, the FSM actively deflects the beam to compensate for angular jitter-induced misalignment between the RX beam and the TX beam. The propagation of the deflected beam through the polarizing optics may cause a change in the polarization state of the TX beam. Fig. 11 illustrates the optical detection scheme for the on-axis LRI. This polarization-dependent detection scheme, comprising PBS1, PBS2, QWP, and QPRs, converts deviations of the TX beam’s polarization state from its nominal configuration (S polarization in this experiment) into a decline in the received RX beam power () at the remote spacecraft, ultimately leading to a reduction in the CNR.
As illustrated in Fig. 11, the LO beam and the RX beam impinge upon the detection system with input powers and , respectively. Propagating the initial Stokes vectors through the optical system via Mueller calculus, we determine the polarization states at QPR1 and QPR2. The corresponding received powers are subsequently computed from the first Stokes parameter . Leveraging the condition characteristic of inter-spacecraft interferometric links [Muller2017PhD], the CNR is derived as
where is the heterodyne efficiency, is the responsivity of the photodiode, is the normalized second Stokes parameter describing the difference between P-pol and S-pol component in a beam, is the electron charge, denotes the power spectral density (PSD) of the photoreceiver’s current, and denotes the PSD of the LO beam power fluctuation normalized to the average power. More details about the derivation of the CNR within the Mueller calculus framework are provided in Appendix B.
During the pointing measurements described in Section IV.2, the polarization state of the TX beam was monitored by the polarimeter on the transponder bench, which recorded the Stokes parameters, azimuth angle, and ellipticity angle. In operational inter-spacecraft links, the TX beam propagates to the remote spacecraft and becomes the RX beam of the optical bench. Based on the polarization state measurements, we analyzed the impact of polarization state variations on the CNR using the relation of Eq. 2.
@@
Fig. 12 shows the variation of the polarization state of the TX beam over 15 hours of continuous measurement. Complementing the pointing measurements described in Section IV.2 (which incorporate hexapod motion), stationary hexapod measurements were taken as well for comparison. For all measurements, the degree of polarization (DoP) of the TX beam remained unity, indicating that depolarization did not occur during the measurements. The DoPs are quantified by the magnitude of the measured normalized Stokes vector: , where denotes the normalized Stokes parameter. The azimuth angles remain confined to 90 while ellipticity angles () approach zero ( 0.11∘), confirming the TX beam’s predominantly S-pol state. During stationary operation, the TX beam exhibited slow variation in Stokes parameters, azimuth angle, and ellipticity angle, which results from the ambient temperature fluctuations. Whereas the hexapod motion and the FSM beam steering produce rapid polarization-state variation, as validated in the pointing measurement results in Fig. 12. During the pointing measurement, the Stokes parameter S1, azimuth angle, and ellipticity angle exhibited PTV fluctuations of 0.0027, 0.89∘, and 1.3∘, respectively. These variations correspond to a 0.14% CNR reduction calculated via Eq. 2. Accounting for the contribution from the ambient temperature variation, the pure beam deflection contribution to CNR loss is expected to be less than the calculated values.
@@ IV.4 Tilt-to-length Coupling
The hexapod was programmed to execute 61-cycle angular scans in the yaw and pitch directions at frequencies of 0.49381 Hz and 0.38532 Hz, respectively. The rotational dynamics of the hexapod conform to a Jerk-limited SGamma profile [NewportHXPManual], which can be mathematically decomposed into a harmonic series within the fundamental mode of the scanning frequency. A detailed elaboration of this motion is provided in Appendix C. The sampling frequency of phase measurement is 101.725 Hz, a multiple of the scanning frequency. During the measurements, the FSM’s strain gauge sensors resolved angular displacements in yaw and pitch with a resolution of 0.05 µrad. In conjunction with the imaging system’s magnification, the FSM’s angular displacements were converted to the rotation angles of the setup, which were then used to determine the TTL coupling.
@@
The linear TTL coupling factor approximates the positional displacement between the rotating pivot point and the CoM [wegener2020tilt, wegener2022analysis]. By leveraging this correlation, the rotating pivot point can be adjusted according to the measured linear TTL coupling factor. The optimal pivot point was determined through iterative measurements until the positional mismatch was reduced below the hexapod’s resolution threshold. The determined optimal pivot point was implemented in the final TTL measurements.
The longitudinal path length variations and TTL coupling with 500 µrad rotation angle in yaw and pitch directions are shown in Fig. 13 (a) and (b), respectively. The QPR1 measurements closely match the QPR2 results, demonstrating the effectiveness of the balanced detection scheme as a redundant design. The longitudinal path length variations in the yaw and pitch directions are constrained to magnitudes below 27 nm and 17 nm, respectively. Correspondingly, the measured TTL coupling values for these directions are less than 370 µm/rad and 350 µm/rad, respectively, exceeding the simulated results by roughly one order of magnitude [yang2022axis].
@@ V SUMMARY AND CONCLUSION
The calibration factors linking the DWS and DPS signals to the tilt angle were determined experimentally based on the dedicated hexapod scans. These calibrated factors were subsequently employed to assess the pointing stability of the system. Under angular jitter analogous to that of the GRACE-FO satellite, the measured pointing stability remained below 10 µrad/ across the frequency range of 2 mHz to 0.5 Hz in both horizontal and vertical directions, meeting the GRACE-FO mission requirements. Additionally, the fluctuation of the TX beam’s polarization state only contributes to a 0.14% CNR degradation during the 15-hour pointing measurements. The TTL coupling was characterized in the laboratory over a tilt angle range of 500 µrad. The measured TTL values exceeded theoretical predictions, with the dominant limitation arising from positioning inaccuracies of the hexapod.
The experimental results validate the principal feasibility of the proposed interferometer and highlight its potential as an alternative for future gravity recovery missions. To achieve higher precision in TTL measurements, it is essential to eliminate longitudinal coupling effects, enhance the hexapod’s positioning accuracy, and improve the optical system’s robustness. Due to the limitations of the breadboard model, phase sensitivity remains unaddressed in this work. Future experimental investigations of the phase noise of the on-axis LRI require characterization in a vacuum chamber using a robust engineering model.
@@ Acknowledgements.
The DWS signals represent the phase differences between beams across the four segments of a QPR, enabling the measurement of relative wavefront tilt between two optical beams [morrison1994automatic, wanner2012methods]. By detecting the phases , , , and at the four segments (A, B, C, and D) of a QPR, the horizontal DWS signal (difference between left and right) and the vertical DWS signal (difference between top and bottom) can be calculated as
Each segment’s phase is determined using an ADPLL [gerberding2013phasemeter, shaddock2006overview].
@@ Appendix B Polarization
Given that in an inter-spacecraft interferometric link, the CNR can be approximated by the expression derived in [Muller2017PhD]:
where is the heterodyne efficiency, is the responsivity of the photodiode, is the electron charge, denotes the power spectral density (PSD) of the photoreceiver’s current, and denotes the PSD of the received LO beam power fluctuation normalized to the average power. The three terms in the denominator (, , and) of Eq. 8 represent the noise of the photoreceiver, shot noise, and relative intensity noise of the laser, respectively. By substituting Eq.5 and Eq.7 into Eq. 8, we derive the polarization-dependent CNR for the interferometric link as
@@ Appendix C Tilt-to-length Coupling
The hexapod is employed to generate periodic rotational motion for TTL measurements. Its rotational dynamics adhere to a jerk-limited SGamma profile, characterized by four parameters: maximum velocity, maximum acceleration, maximum jerk time, and minimum jerk time [NewportHXPManual]. The rotation angle of the hexapod can be decomposed into a series of harmonics. Specifically, with the fundamental mode , the rotation angle at time can be expressed as
@@
During the TTL measurement, the hexapod executed 61 cycles of angular scans in the yaw and pitch directions, respectively. With the beam steering loops closed, the FSM’s angular displacements, measured via strain gauge sensors, were converted to equivalent hexapod rotation angles by multiplying the angular magnification of the imaging system. Using the measured angles, the resulting angular displacement spectral densities, computed via logarithmic frequency axis power spectral density (LPSD) [trobs2006improved], are presented in Fig. 14. Notably, our rotational scans revealed a series of harmonics in addition to the predominant fundamental mode.

### v2

@@ IV.1 Angle calibration
Residual alignment imperfections in the QPR introduce cross-coupling effects in the measurements. As shown in Fig. 9 (c), (d), (e), and (f), jumps in cross-coupling signals are observed when the hexapod reverses its scanning direction, while the cross-coupling signals remain relatively constant between these jumps. These jumps may result from the hysteresis and backlash effects [24], which produce transient instability of the hexapod’s motion during its settling time. Excluding the observed jumps, a linear fit of the measured cross-coupling DWS and DPS signals with respect to the rotation angle determines the cross-coupling factors ( and ), as presented in Table 1.
@@ IV.2 Beam Pointing
The stability of the beam pointing in an LRI system affects the laser link, and the pointing noise can be coupled into ranging measurements through TTL coupling. To investigate the pointing stability of the on-axis architecture, the hexapod was commanded to perform typical motions derived from GRACE-FO pointing data recorded on January 1st, 2019 [9], which artificially produce the TX beam pointing errors. The hexapod motion was updated at a frequency of 2 Hz over 15-hour continuous measurements, with the CoM as the nominal pivot point. The beam steering loops were closed during the hexapod motions. The coordinate information, provided from the hexapod’s position sensors, was used to monitor hexapod motion in real time. Accordingly, the resulting amplitude spectral density (ASD) of the hexapod motion is shown in Fig. 10 (a), which is consistent with the result in [9]. Although several harmonics above 0.1 Hz are present, these frequency components are at the boundary of the GRACE-FO frequency band of 2 mHz to 0.1 Hz [3] and thus do not affect our analysis here.
To evaluate the co-alignment between the RX beam and the TX beam, the DWS signals from QPR1 and QPR2 were recorded. For comparison, the DPS signals from QPR3 were used to characterize the absolute TX beam pointing. Using the relation of Eq. 1, the pointing angles were determined by multiplying the DWS and DPS signals by the inverse of the corresponding conversion matrices detailed in Table 1. The corresponding ASD of the pointing angle is displayed in Fig. 10 (b). In addition to the in-loop DWS measurements from the DWS1, out-of-loop DWS measurements from the DWS2 were performed as well. The horizontal (yaw) and vertical (pitch) co-alignments derived from the in-loop DWS measurements are below 10-10 rad/ at the frequency range from 0.07 mHz to 0.5 Hz, which indicates that the error signals are effectively suppressed by the control loops. For frequencies between 0.2 mHz and 0.1 Hz, both direction co-alignments derived from the out-of-loop DWS measurements are lower than 10 µrad/, which fulfills the pointing requirement of NGGM mission [25]. Compared to the pointing measurements of the GRACE-FO LRI [29], our measurements exhibit better pointing stability in the frequency range below 0.1 Hz.
@@
Following the activation of the beam steering loops, the beam pointing stability in the yaw and pitch directions (Fig. 10 (a)) was significantly enhanced, reaching the performance levels indicated by the orange and gray lines in Fig. 10 (b). This represents a substantial improvement, particularly in the frequency range below 0.01 Hz where a two-order-of-magnitude reduction is observed. As the critical component in the beam steering loops, the noise performance of the FSM impacts the beam pointing stability. In the DWS2 measurements, thermal and air fluctuations limit further improvement at low frequencies. Additionally, although the imaging system maintains the optical path length almost consistent, the hexapod motion and the steering loops introduce variations in the laser propagation path on the optical bench, further constraining the co‑alignment of the two beams in the low-frequency band.
@@ IV.3 Polarization
In the beam steering loop, the FSM actively deflects the beam to compensate for angular jitter-induced misalignment between the RX beam and the TX beam. The propagation of the deflected beam through the polarizing optics may cause a change in the polarization state of the TX beam. Fig. 11 illustrates the optical detection scheme for the on-axis LRI. This polarization-dependent detection scheme, comprising PBS1, PBS2, QWP, and QPRs, converts deviations of the TX beam’s polarization state from its nominal configuration (S polarization in this experiment) into a decline in the received RX beam power () at the remote spacecraft, ultimately leading to a reduction in the .
As illustrated in Fig. 11, the LO beam and the RX beam impinge upon the detection system with input powers and , respectively. Propagating the initial Stokes vectors through the optical system via Mueller calculus, we determine the polarization states at QPR1 and QPR2. The corresponding received powers are subsequently computed from the first Stokes parameter . Leveraging the condition characteristic of inter-spacecraft interferometric links [23], the is derived as
where is the heterodyne efficiency, is the responsivity of the photodiode, is the normalized second Stokes parameter describing the difference between P-pol and S-pol component in a beam, is the electron charge, denotes the power spectral density (PSD) of the photoreceiver’s current, and denotes the PSD of the LO beam power fluctuation normalized to the average power. More details about the derivation of the within the Mueller calculus framework are provided in Appendix B.
During the pointing measurements described in Section IV.2, the polarization state of the TX beam was monitored by the polarimeter on the transponder bench, which recorded the Stokes parameters, azimuth angle, and ellipticity angle. In operational inter-spacecraft links, the TX beam propagates to the remote spacecraft and becomes the RX beam of the optical bench. Based on the polarization state measurements, we analyzed the impact of polarization state variations on the using the relation of Eq. 2.
@@
Fig. 12 shows the variation of the polarization state of the TX beam over 15 hours of continuous measurement. Complementing the pointing measurements described in Section IV.2 (which incorporate hexapod motion), stationary hexapod measurements were taken as well for comparison. For all measurements, the degree of polarization (DoP) of the TX beam remained unity, indicating that depolarization did not occur during the measurements. The DoPs are quantified by the magnitude of the measured normalized Stokes vector: , where denotes the normalized Stokes parameter. The azimuth angles remain confined to 90 while ellipticity angles () approach zero ( 0.11∘), confirming the TX beam’s predominantly S-pol state. During stationary operation, the TX beam exhibited slow variation in Stokes parameters, azimuth angle, and ellipticity angle, which results from the ambient temperature fluctuations. Whereas the hexapod motion and the FSM beam steering produce rapid polarization-state variation, as validated in the pointing measurement results in Fig. 12. During the pointing measurement, the Stokes parameter S1, azimuth angle, and ellipticity angle exhibited PTV fluctuations of 0.0027, 0.89∘, and 1.3∘, respectively. These variations correspond to a 0.14% reduction in the calculated via Eq. 2. Accounting for the contribution from the ambient temperature variation, the pure beam deflection contribution to loss is expected to be less than the calculated values.
@@ IV.4 Tilt-to-length Coupling
The hexapod was programmed to execute 61-cycle angular scans in the yaw and pitch directions at frequencies of 0.49381 Hz and 0.38532 Hz, respectively. The rotational dynamics of the hexapod conform to a Jerk-limited SGamma profile [24], which can be mathematically decomposed into a harmonic series within the fundamental mode of the scanning frequency. A detailed elaboration of this motion is provided in Appendix C. The sampling frequency of phase measurement is 101.725 Hz, a multiple of the scanning frequency. During the measurements, the FSM’s strain gauge sensors resolved angular displacements in yaw and pitch with a resolution of 0.05 µrad. In conjunction with the imaging system’s magnification, the FSM’s angular displacements were converted to the rotation angles of the setup, which were then used to determine the TTL coupling factor.
@@
The linear TTL coupling factor approximates the positional displacement between the rotating pivot point and the CoM [40, 41]. By leveraging this correlation, the rotating pivot point can be adjusted according to the measured linear TTL coupling factor. The optimal pivot point was determined through iterative measurements until the positional mismatch was reduced below the hexapod’s resolution threshold. The determined optimal pivot point was implemented in the final TTL measurements.
The longitudinal path length variations and TTL coupling factor with 500 µrad rotation angle in yaw and pitch directions are shown in Fig. 13 (a) and (b), respectively. The QPR1 measurements closely match the QPR2 results, demonstrating the effectiveness of the balanced detection scheme as a redundant design. The longitudinal path length variations in the yaw and pitch directions are constrained to magnitudes below 27 nm and 17 nm, respectively. Correspondingly, the measured TTL coupling values for these directions are less than 370 µm/rad and 350 µm/rad, respectively, exceeding the simulated results by roughly one order of magnitude [43]. The GRACE-FO LRI requires an optical bench TTL coupling of less than 80 µm/rad within a range of 300 µrad [26], with similar requirements for the GRACE-C and NGGM missions. However, the measured TTL numbers within this range are 145 µm/rad (yaw) and 188 µm/rad (pitch), which remain approximately twice the specified limit.
@@
The longitudinal coupling can be estimated using Eq. 12. During 500 µrad rotations, pivot point position uncertainties of 4 µm in directions and 2 µm in direction(as specified in [24]) resulted in measurement uncertainty of 9.74 nm and 4.0 nm in the longitudinal pathlength variation. These correspond to 36.1% and 23.5% of the measured magnitudes in yaw and pitch direction, respectively. Since the TTL coupling factor was determined by the difference between the measured longitudinal path length at adjacent angles, such significant uncertainty inevitably leads to substantial errors in the TTL coupling factor, as shown in Fig. 13(b). Further details regarding the estimation methodology are provided in the Appendix C.
@@ V SUMMARY AND CONCLUSION
The calibration factors linking the DWS and DPS signals to the tilt angle were determined experimentally based on the dedicated hexapod scans. These calibrated factors were subsequently employed to assess the pointing stability of the system. Under angular jitter analogous to that of the GRACE-FO satellite, the measured pointing stability remained below 10 µrad/ across the frequency range of 0.2 mHz to 0.5 Hz in both horizontal and vertical directions, meeting the GRACE-FO mission requirements. Additionally, the fluctuation of the TX beam’s polarization state only contributes to a 0.14% degradation during the 15-hour pointing measurements. The TTL coupling was characterized in the laboratory over a tilt angle range of 500 µrad. The measured TTL values exceeded theoretical predictions, with the dominant limitation arising from positioning inaccuracies of the hexapod.
The experimental results validate the principal feasibility of the proposed interferometer and highlight its potential as an alternative for future gravity recovery missions. The concepts and technologies developed herein can benefit future inter-spacecraft laser interferometry missions. To achieve higher precision in TTL measurements, it is essential to eliminate longitudinal coupling effects, enhance the hexapod’s positioning accuracy, and improve the optical system’s robustness. Due to the limitations of the breadboard model, phase sensitivity remains unaddressed in this work. Future experimental investigations of the phase noise of the on-axis LRI require characterization in a vacuum chamber using a robust engineering model.
@@ Acknowledgements.
The DWS signals represent the phase differences between beams across the four segments of a QPR, enabling the measurement of relative wavefront tilt between two optical beams [22, 38]. By detecting the phases , , , and at the four segments (A, B, C, and D) of a QPR, the horizontal DWS signal (difference between left and right) and the vertical DWS signal (difference between top and bottom) can be calculated as
@@ Appendix B Polarization
Given that in an inter-spacecraft interferometric link, the can be approximated by the expression derived in [23]:
where is the heterodyne efficiency, is the responsivity of the photodiode, is the electron charge, denotes the power spectral density (PSD) of the photoreceiver’s current, and denotes the PSD of the received LO beam power fluctuation normalized to the average power. The three terms in the denominator (, , and) of Eq. 8 represent the noise of the photoreceiver, shot noise, and relative intensity noise of the laser, respectively. By substituting Eq.5 and Eq.7 into Eq. 8, we derive the polarization-dependent for the interferometric link as
@@ Appendix C Tilt-to-length Coupling
The hexapod is employed to generate periodic rotational motion for TTL measurements. Its rotational dynamics adhere to a jerk-limited SGamma profile, characterized by four parameters: maximum velocity, maximum acceleration, maximum jerk time, and minimum jerk time [24]. The rotation angle of the hexapod can be decomposed into a series of harmonics. Specifically, with the fundamental mode , the rotation angle at time can be expressed as
@@
During the TTL measurement, the hexapod executed 61 cycles of angular scans in the yaw and pitch directions, respectively. With the beam steering loops closed, the FSM’s angular displacements, measured via strain gauge sensors, were converted to equivalent hexapod rotation angles by multiplying the angular magnification of the imaging system. Using the measured angles, the resulting angular displacement spectral densities, computed via logarithmic frequency axis power spectral density (LPSD) [36], are presented in Fig. 14. Notably, our rotational scans revealed a series of harmonics in addition to the predominant fundamental mode.
Similar to the off-axis LRI, the round-trip longitudinal pathlength variation that arises from the offset between the pivot point and the reference point during the rotation is given by
where defines the rotation matrix, and denotes the unit vector along the line-of-sight direction (the laser propagation direction from the reference bench to the transponder bench) [41]. For small rotation angles and , a first‑order approximation yields
with and denoting the mismatches in the hexapod’s rotation coordinate frame [28]. Since the rotation frame used in the TTL measurements was derived from the hexapod’s default coordinate system via a 14.4∘ yaw rotation transformation, this additional transfer matrix must be considered when estimating the longitudinal coupling. By applying Eq. 12, the positional errors from the specification and rotation angles, the potential longitudinal coupling can be determined.